OpenStudy (anonymous):

Write the slope intercept form of the equation of the line through the given point with the given slope Through: (1,4) Slope =undefined

OpenStudy (jonnyvonny):

An undefined slope is when there is a horizontal line or vertical line. Which would be, in this case, y=4, I believe. It can also be x=1.

OpenStudy (anonymous):

So both could be the form of the equation?

OpenStudy (jonnyvonny):

However, I forgot to mention, since they say slope intercept form, that is y=mx+b. y=(0)x+4 y=4

OpenStudy (anonymous):

Ahhhh ok m acts as the slop and just plug in x and y

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Wait why is b 4 ?

OpenStudy (jonnyvonny):

Indeed, the m will always be 0, since its undefined. They give you the (1,4), so plug in the points.

OpenStudy (anonymous):

Oh nevermind b is the y intercept, allright thanks for the help

OpenStudy (jonnyvonny):

4=mx+b 4=0(1)+b 4=b From then, we can write a general equation that applies to all points; y=(0)x+4 y=4